OpenStudy (anonymous):

I need help this project is hard and no way to reach my teachers please help

OpenStudy (anonymous):

There are many different proofs of the Pythagorean Theorem. Researchand choose three different proofs. Outline and provide an explanation foreach. Be sure to site the sources of information you used.

OpenStudy (anonymous):

http://www.cut-the-knot.org/pythagoras/index.shtml

OpenStudy (anonymous):

I got proofs I just need help with outline and explanation

OpenStudy (anonymous):

I think those are good explanations

OpenStudy (anonymous):

@ArkGoLucky I got the sites I need help with outline and explanation that's hard

OpenStudy (anonymous):

can you post the proofs that you have?

OpenStudy (anonymous):

how do you outline it

OpenStudy (anonymous):

Proof #3 already well so far I got stuck at outline so I choose one for now.

OpenStudy (anonymous):

there's not really a standard for outlining proofs. I think that's sort of teacher dependent. I would assume by outline your teacher means to list the steps of the proof

OpenStudy (anonymous):

so how would you outline proof 3 then

OpenStudy (anonymous):

1. Draw four identical right triangles all rotated 0 degrees, 90 degrees, 180 degrees and 270 degrees. 2. Label the corresponding sides of the triangle a, b and c. 3. Without rotating any triangle, put the triangles in a square of side length c, which should be the hypothenuse of all the triangles. 4. The square has a square in the middle with a side length of (a-b) 5. The area of all the triangles is 4*(ab/2) or 2ab. 6. The area of the square is c^2 and also the sum of area of the triangles and the square in the middle so c^2=(a-b)^2 + 2ab 7. Expanding the right side of the equation, we get c^2=a^2+b^2